OpenStudy (lidia.12):

Let f(x) =3x, g(x)=x+1. Find the composite function. (g•f)(2) PreCalculus

OpenStudy (lowkey.s):

I really don't mate. sorry

OpenStudy (mhchen):

I don't really know either, sorry. Just kidding I do know how to help you. Any ideas?

jimthompson5910 (jim_thompson5910):

Hints: \[\Large (g \circ f)(x) = g(f(x))\] \[\Large (g \circ f)(2) = g(f(2))\]

OpenStudy (mhchen):

Second of all, the composite of a function like g o f is g(f(x))

OpenStudy (mhchen):

So you find the value of f(2), then you PLUG that value into g(value) So f(2) = 3(2) = 6 g(6) = 6 + 1 <-- Solve this And that's how you do composite functions

OpenStudy (lidia.12):

What about the same thing but (f•g)(2)? Is it basically the same thing?

OpenStudy (agent0smith):

I'm pretty sure this question is multiplying functions...? Might want to clarify with a screenshot.

jimthompson5910 (jim_thompson5910):

If you mean \[\Large (f \circ g)(2)\] Then \[\Large (f \circ g)(2) = f(g(2))\]